OpenStudy (anonymous):

find the exact value by using half angle identities. cos165°

OpenStudy (anonymous):

1 - cos(330°) ------------------- = sin² 165° ......2 1 - √(3)/2 ------------- = sin² 165° ......2 .0670 = sin² 165° .2588 = sin 165° <== ANSWER Remember that sine values in the second quadrant are positive. 1 + cos(330°) ------------------- = cos² 165° ......2 1 + √(3)/2 -------------- = cos² 165° ......2 .9330 = cos² 165° -.9659 = cos 165° <== ANSWER

OpenStudy (anonymous):

that fine?

OpenStudy (dumbcow):

 from this, you can say \[\cos(165) = -\cos(15) = -\cos(\frac{30}{2})\]

OpenStudy (anonymous):

Thanks to both! :D